OpenStudy (anonymous):

you reach into a bag of coins and withdraw two coins. what is the probability you withdrew a nickel and then a dime if the bag held five pennies, ten nickles, and four dimes?

OpenStudy (anonymous):

@Haichi ?

OpenStudy (anonymous):

\[\frac{10}{19}\times \frac{4}{18}\]

OpenStudy (anonymous):

Alright, what do I do with those two fractions? O.o

OpenStudy (anonymous):

Is the answer 40/361?